# Personalized Generative Low-light Image Denoising and Enhancement

### v2

@@ 4.1 Dataset
4.2 Implementation Details
Our denoising model is built on the ADM architecture [12], and the ID-consistent physical extractor is implemented using the aggregation network from LAP [65]. For training, we use a P2-weighted loss function [8], which measures the discrepancy between the predicted and ground-truth noise. This loss is defined as: $L=\lambda_{t}||\hat{\epsilon}_{t}-\epsilon_{t}||_{2}^{2}$ , where $\lambda_{t}$ serves as a hyperparameter to adjust the loss weight dynamically across different timesteps.
@@ 4.2 Implementation Details
During the training stage, we train our model on the CelebAMask-HQ dataset [30], comprising 30,000 images distributed across 4,516 unique identities(IDs). We use 3318 IDs for training and 200 IDs for simulated testing. Within each ID, there are 3–20 images. The number of gallery images we used per person varies from 3 to 6, with most around 5. The Adam optimizer is applied with a learning rate of 1e-4. Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
4.3 Performance in Real-Captured Scenarios
@@
Testing model performance in real-life low-light scenarios is crucial and challenging, as real-world conditions often involve uncontrollable factors compared to simulated ones. Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios , as illustrated in Section 4.1.
@@
We compare our methods with several state-of-the-art works: UTV-Net[66], GFPGAN[52], SNR-aware[59], MIRNet[61, 62], DiffLL[22], Codeformer [67]. These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research. Particularly, GFPGAN and Codeformer are designed specifically for face image restoration, which is highly competitive with our method. We re-trained/fine-tuned all the baselines with the same training dataset as ours for fair comparison.
@@
Quantitative and qualitative results are presented. To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [20] and the Kernel Inception Distance (KID) [1], as these metrics offer a robust evaluation aligned with human perception. Since ground-truth images are unavailable for real-captured cases, we compute the FID and KID score between the model’s output and the corresponding subjects’ gallery photos. As shown in Fig. 6, our method produces images with the highest quality and level of detail. The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips, while maintaining the highest consistency in facial identity.
@@
Since our work focuses on face images, we also check on the identity preservation performance. Besides the image quality metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[11] to further evaluate the identity consistency. Again, since ground truth images are unavailable for real-captured cases, we calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo of the corresponding subjects. The results are also shown in Table 2. Again, our method has the highest identity score, and the visual results are consistent with the quantitative metric results. The results demonstrate the supreme ability of our proposed DiffPGD model on real low light denoising and enhancement task.
4.4 Performance in Simulated Scenarios
In simulated scenarios, We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set. We simulated various low-light noisy levels by setting different photons per pixel (PPP), i.e., the average number of photons arriving at the sensor within the exposure time. The lower the PPP, severe the degradation is. A lower photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
Quantitative results are provided in Table 1. The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best. (Refer to supplementary for more visual comparisons). From the table, we could see that CodeFormer surpasses us on PSNR. Fig. 7 illustrates a visual comparison between our method and CodeFormer under inputs ranging from 5 to 65 PPP. We observe that CodeFormer tends to generate smoother images while we could preserve better high-frenquency details. Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 PPP to 13 PPP, while exhibiting fewer artifacts. The advantage of our model in maintaining identity consistency is particularly evident in heavily degraded cases. The results highlight the strong performance of our proposed DiffPGD model in low-light denoising and enhancement tasks.
@@ 4.4 Performance in Simulated Scenarios
SNR-aware[55] 252.2397 0.1264 0.1990
UTVNet[66] 341.8502 0.1617 0.2002
MIRNet[61, 62] 276.3853 0.1659 0.2210
DiffLL[22] 252.1553 0.1421 0.3196
GFPGAN[52] 252.7411 0.1249 0.1525
CodeFormer[67] 281.9744 0.1459 0.2965
Ours 251.9318 0.0882 0.3708
Table 2: Comparisons with state-of-the-art works on the real-captured test cases.
@@
w/o physical buffers 272.6148 0.1281 0.3169
w/ physical buffers 251.9318 0.0882 0.3708
Table 3: Ablation comparisons of using ID physical buffers on real-captured photo test.
Figure 7: Visual comparison on the simulated cases. Our method maintains good identity consistency even at extremely low PPP and exhibits fewer artifacts.
@@
5 Ablation Study
ID-Consistent Physical Buffers for Identity Preservation. We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-captured testing scenarios. In Table 3, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases. The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
Effect of the Number of Gallery Images. The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos. As claimed in the paper, facial features extracted from a single image are inherently limited. In Fig. 8, we evaluate the effect of the number of gallery images used to generate the ID physical buffer. We analyze it with the real-captured set where 6 gallery images were collected from each subject. The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image. Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case without gallery photos.
6 Conclusion
In this work, we addressed the challenges of low-light facial image restoration by introducing Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that leverages user-specific photo galleries to improve denoising performance. By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning. Our experiments, which included extensive evaluations on multiple challenging datasets, demonstrated that DiffPGD not only outperforms existing diffusion-based denoising methods but also delivers superior quantitative and qualitative results across various testing scenarios. This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving the way for personalized and more robust image restoration techniques.
Supplementary Material
Figure 9: Simulation process for generating low-light noisy face images.
Appendix A Dataset
As mentioned in Section $4.1$ of the main paper, since no open-sourced low-light noisy datasets are available, we adopt a simulation approach to generate low-light noisy face images for training our model. Specifically, we utilize the CelebAMask-HQ dataset’s [30] face images as the ground truth RGB images.
As shown in Fig. 9, the first step in the simulation process involves passing the RGB image through a pre-trained Inverse ISP network [34] to generate a pseudo ground truth RAW mosaic image. For visualization purposes, we display the demosaiced version of this image. We utilize the pretrained weights provided by the authors [34], which were obtained by training the network on iPhone camera images, for unprocessing the RGB images into RAW format.
This pseudo ground truth RAW image is then passed through a Poisson-Gaussian Simulator to simulate the low-light noisy RAW mosaic image. It is noteworthy that during the training process, we vary the camera sensor parameters (in the Poisson-Gaussian simulator) across a wide range of values to enable the model to handle diverse camera sensor noise profiles. The simulation setup assumes a $12$-bit image sensor with parameters similar to those of Canon DSLR cameras. The parameter ranges are summarized in Table 5.
Table 5: Simulation parameters used in the Poisson-Gaussian Simulator.
  Parameter Range
Full Well Capacity $[19000\text{e}^{-},64000\text{e}^{-}]$
Quantum Efficiency $[0.32,0.54]$
Dark Current $[2.2\text{e}^{-},11.7\text{e}^{-}]$
Read Noise $[2.2\text{e}^{-},10.8\text{e}^{-}]$
Photons-per-Pixel (PPP) $[13\,\text{PPP},65\,\text{PPP}]$
These parameters are carefully chosen to ensure realistic noise characteristics during training.
Appendix B Additional Visual Results
In Fig. 12, Fig. 14, and Fig.15, we present additional visual comparisons between our model and baseline methods on real-captured and simulated cases. Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
Figure 10: A network architecture diagram for identity feature restoration using a face encoder.
  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
w/ Encoder 12.62 0.2574 0.6315 0.6906
w/ physical buffers 13.85 0.3753 0.6042 0.8039
@@ Appendix B Additional Visual Results
Table 6: Comparison of ID physical buffers and encoder-based methods on the real-captured photo test set. For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
Appendix C More Implementation Details
During training and testing, the image size is 256 $\times$ 256. The architecture of the diffusion model is adapted from [12]. We modify the architecture so that the model can take the concatenation of physical buffers and degraded images as conditions. The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000, and the two hyperparameters P2_gamma and p2_k are both set to 1.0. For additional architectural details, refer to [12]. During inference, we adopt the DDIM method [46] with 100 denoising steps. The model is trained using the Adam optimizer [27] with a learning rate of $10^{-4}$. The weight decay is set to 0. All experiments are performed on a single NVIDIA A100 GPU.
Appendix D Encoder vs. Physical Buffer?
Many previous studies have used encoder-based methods to extract latent facial feature information from reference photos[51]. To evaluate whether the encoder can effectively extract enough facial features to guide the diffusion model in solving the low-light enhancement and denoising task, we conducted an experiment. As illustrated in Fig. 10, we utilized an encoder (ResNet-18 [19]) to extract latent codes from the target individual’s gallery photos. Following the method employed by ADM [12] for time embedding, we scaled and shifted the features in each layer of the diffusion model using the latent codes. This process integrates the identity information extracted by the encoder into the diffusion model.
@@ Appendix C More Implementation Details
We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set. This is a special semi-real set, in which we take photos of some printed paper photos at night. We prepare this semi-real set because The ground-truth will be available to be compared with. The quantitative and visual comparisons are presented in Table 6 and Fig. 11, respectively. The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach. One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery. In contrast, the encoder-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle. Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
Appendix E Gallery Photos
Fig. 13 presents the the gallery photos for the test subjects included in the Fig. 6 of the main paper.
Figure 11: Visual comparison on different identity representation strategies (physical buffer and face encoder).
Figure 12: Additional real-captured visual comparison.
Figure 13: Gallery photos.
Figure 14: Additional visual comparison.
Figure 15: Additional visual comparison.

### v3

@@ IV-A Dataset
Figure 7: Simulation process for generating low-light noisy face images. A ground truth RGB image is first passed through a pre-trained Inverse ISP [li2024efficient] to obtain a pseudo ground truth RAW mosaic. This is followed by applying a Poisson-Gaussian noise model with varying sensor parameters to simulate a low-light noisy RAW. Finally, a Forward ISP module converts the noisy RAW back to a low-light noisy RGB image.
IV-B Implementation Details
Our denoising model is built upon the ADM architecture [dhariwal2021diffusion]. As illustrated in Fig. 6, we modify the original architecture so that the model can take concatenated degraded images as conditional inputs. In addition, we employ an encoder (ResNet-18 [he2016deep]) to extract a global latent code from the target individual’s physical buffers, which are obtained from the gallery photos. Following the FiLM-style [Perez2017FiLMVR] modulation used for time embeddings in ADM [dhariwal2021diffusion], this latent code modulates the UNet features in each residual block via scale–and–shift or additive transforms, thereby injecting identity information into the diffusion model. The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000. For additional architectural details, refer to [dhariwal2021diffusion].
The ID-consistent physical extractor is directly initialized from the pre-trained weights of the aggregation network in LAP [Zhang2021lap], without the need of retraining.
@@ IV-B Implementation Details
For training our model, we use the loss function defined in Eq. (10). The image size is 256 $\times$ 256. During the training stage, we train our model on the CelebAMask-HQ dataset [Lee2020CelebAMaskhq], comprising 30,000 images distributed across 4,516 unique identities (IDs). We use 3318 IDs for training and 200 IDs for simulated testing. Within each ID, there are 3–20 images. The number of gallery images we used per person varies from 3 to 6, with most of them around 5. During inference, we used 3 to 6 images per person for the simulated test set, and 6 gallery images per person for the real-captured test set.
The model is trained using the Adam optimizer [kingma2014adam] with a learning rate of $10^{-4}$. The weight decay is set to 0. Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model. During inference, we adopt the DDIM method [song2020denoising] with 200 denoising steps. All experiments are performed on a single NVIDIA A100 GPU.
IV-C Performance in Real-Captured Scenarios
@@
Testing model performance in real-life low-light scenarios is crucial and challenging, as real-world conditions often involve uncontrollable factors compared to simulated ones. Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios, our procedure is described in Section IV-A.
@@
We compare our method with several state-of-the-art methods: UTV-Net [zheng2021adaptive], GFPGAN[wang2021towards], SNR-aware[xu2022snr], MIRNet[zamir2020learning], DiffLL[jiang2023low], Codeformer [zhou2022towards]. These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research. Particularly, GFPGAN[wang2021towards] and Codeformer[zhou2022towards] are designed specifically for face image restoration, which is highly competitive with our method. We re-trained/fine-tuned all the baselines with the same training dataset as ours for fair comparison.
@@
Quantitative and qualitative results are presented. To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [Heusel2017FID] and the Kernel Inception Distance (KID) [Binkowski2018KID], as these metrics offer a robust evaluation aligned with human perception. Since ground-truth images are unavailable for real-captured cases, we compute FID and KID scores between the model’s output and the corresponding subjects’ gallery photos. The results are shown in Table VI. Our method achieves the best FID score and the second-best KID score. When examining the visual results presented in Figs. 8, 9 and 10, our method produces images with the highest perceptual quality and level of detail. The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips.
Figure 8: Visual comparison on real cases. The first image presents the input images along with their capture parameters. In the second image, we enhance the brightness to better visualize the noise present in the input. Please zoom in for the best visual experience.
Figure 9: Visual comparison on real cases. The first image presents the input images along with their capture parameters. In the second image, we enhance the brightness to better reveal the noise in the input. Please zoom in for the best visual experience.
Figure 10: Visual comparison on real cases. The first image presents the input images along with their capture parameters. In the second image, we enhance the brightness to better reveal the noise inside the input. Please zoom in for the best visual experience.
Figure 11: Visual comparison on the simulated cases. The lower the ppp, the more severe the degradation is. Our method maintains good identity consistency even at extremely low ppp and exhibits fewer artifacts. Please zoom in for the best visual experience.
@@
Since our work focuses on face images, we also check on the identity preservation performance. Besides the image quality metrics, we adopt the identity score, which computes the cosine similarity of face embeddings extracted by the FaceNet model [Schroff2015FaceNetAU] pretrained on the VGGFace2 dataset [Cao2017VGGFace2AD], to evaluate identity consistency. Again, since ground truth images are unavailable for real-captured cases, we calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo of the corresponding subjects. The results are shown in Table VI. Our method has the highest identity score, and the visual results are consistent with the quantitative metric results, maintaining the highest consistency in facial identity. The results demonstrate the ability of our proposed DiffPGD model on real low light denoising and enhancement tasks.
TABLE VI: Comparisons with state-of-the-art works on the real-captured test cases.
@@
SNR-aware[xu2022snr] 358.03 0.3206 0.56
UTVNet[zheng2021adaptive] 360.13 0.2304 0.42
MIRNet[zamir2020learning] 274.57 0.1732 0.63
DiffLL[jiang2023low] 301.13 0.2496 0.69
GFPGAN[wang2021towards] 276.30 0.2023 0.30
CodeFormer[zhou2022towards] 295.68 0.2158 0.61
Ding et al. [ding2024restoration] 295.85 0.2438 0.24
FourierDiff[Lv2024FourierPD] 303.78 0.2383 0.70
Ours 270.14 0.1765 0.72
IV-D Performance in Simulated Scenarios
In simulated scenarios, we use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set. We simulated various low-light noisy levels by setting different photons per pixel (ppp), i.e., the average number of photons arriving at the sensor within the exposure time. A low photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
Quantitative results are provided in Table VII. The image restored by our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistency. From the table, we could see that CodeFormer surpasses us on PSNR value. Fig. 11 illustrates a visual comparison between our method and state-of-the-art works under inputs ranging from 5 to 39 ppp, which corresponds to $\sim$0.049 lux to 0.384 lux, estimated at the sensor side assuming typical low-light CMOS settings with pixel pitch in the range of 2–3 $\mu\text{m}$, quantum efficiency 0.3–0.54, and short integration time on the order of 10–30 ms. We observe that CodeFormer tends to generate smoother images while we could preserve better high-frequency details. Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 ppp to 13 ppp ($\sim$0.049 lux to 0.128 lux), while exhibiting fewer artifacts. The advantage of our model in maintaining identity consistency is particularly evident in heavily degraded cases. The results highlight the strong performance of our proposed DiffPGD model in low-light denoising and enhancement tasks. Like any denoising method, there is a physical limit beyond which our method will break down. This happens when input image is below 5ppp ($\sim$0.049 lux). However, in this extremely low photon condition, we still perform much better than competing methods.
TABLE VII: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5ppp ($\sim$0.049 lux), 10ppp($\sim$0.098 lux), 13ppp($\sim$0.128 lux), 26ppp($\sim$0.256 lux), and 39ppp($\sim$0.384 lux). The lower the ppp, the more severe the degradation is.
  lux 0.049 lux 0.098 lux 0.128 lux 0.256 lux 0.384 lux
Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$
SNR-aware[xu2022snr] 0.3093 200.10 0.1382 0.5580 134.44 0.0664 0.6366 120.44 0.0526 0.7922 96.55 0.0330 0.8547 84.36 0.0244
MIRNet[zamir2020learning] 0.1256 376.98 0.4035 0.3596 320.81 0.2719 0.4698 303.99 0.2506 0.7068 209.01 0.1579 0.7928 161.86 0.1106
FourierDiff[Lv2024FourierPD] 0.0451 384.58 0.4545 0.2045 308.99 0.2753 0.3233 275.99 0.2121 0.6118 202.49 0.1462 0.7195 163.25 0.1137
DiffLL[jiang2023low] 0.3835 190.41 0.1687 0.5496 167.77 0.1390 0.6228 155.87 0.1244 0.7796 128.73 0.0898 0.8388 115.59 0.0776
GFPGAN[wang2021towards] 0.0070 388.93 0.4256 0.1194 233.32 0.1522 0.2851 203.81 0.1152 0.6175 123.27 0.0624 0.7294 106.47 0.0497
CodeFormer[zhou2022towards] 0.3553 111.46 0.0491 0.5867 91.13 0.0326 0.6501 84.80 0.0281 0.7998 80.66 0.0264 0.8645 64.04 0.0167
Ours 0.4527 85.50 0.0178 0.6236 75.39 0.0135 0.6923 75.65 0.0179 0.8168 68.44 0.0114 0.8719 62.73 0.0112
V Ablation Study
ID-Consistent Physical Buffers for Identity Preservation. In Table VIII, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases. The results demonstrate that using physical buffers significantly increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
TABLE VIII: Ablation comparisons of using ID physical buffers on real-captured test cases.
@@ V Ablation Study
w/o physical buffers 281.68 0.1814 0.68
w/ physical buffers 270.14 0.1765 0.72
Effect of the Number of Gallery Images. The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the user’s gallery photos. As stated in the paper, facial features extracted from a single image are inherently limited. In Fig. 12, we evaluate the effect of the number of gallery images used to generate the ID physical buffer. We analyze it with the real-captured set where 6 gallery images were collected from each subject. The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image. Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it.
@@
Direct-Gallery vs. Physical Buffer? Some previous studies have attempted to extract latent facial feature representations directly from reference photos [Varanka2024gallery]. To evaluate whether this strategy can effectively capture sufficient facial information to guide the diffusion model, compared with extraction from physical buffers, we conduct an experiment. As illustrated in Fig. 13, instead of deriving features from physical buffers, we directly extract latent codes from the target individual’s gallery photos.
Figure 13: A network architecture diagram for identity feature restoration. It extracts latent identity features from the user’s photo gallery directly with an encoder, which are then used to condition a denoising diffusion model.
@@
TABLE IX: Comparison between ID physical-buffer-based and direct-gallery-based methods on the real-captured photo test set. For a fair comparison, the same gallery set is used for each individual in both the direct-gallery-based and physical-buffer-based approaches.
  Methods LPIPS $\downarrow$ SSIM $\uparrow$ KID $\downarrow$ ID Score $\uparrow$
w/ direct-gallery 0.6316 0.2574 0.0371 0.77
w/ physical buffers 0.5742 0.4175 0.0161 0.83
Figure 14: Visual comparison on different identity representation strategies.
@@
We compare the restored images generated with the direct-gallery-based method with the proposed physical buffer-guided method on our real-captured photo set. This is a special semi-real set, in which we take photos of printed paper photos at night, allowing ground-truth reference for evaluation. The quantitative and visual comparisons are presented in Table IX and Fig. 14, respectively. The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the direct-gallery-based approach. One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery. In contrast, the direct-gallery-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle. Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
VI Conclusion
In this work, we addressed the challenges of low-light facial image restoration by introducing Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that leverages user-specific photo galleries to improve denoising performance. By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning. Our experiments, which included extensive evaluations on multiple challenging datasets, demonstrated that DiffPGD not only outperforms existing diffusion-based denoising methods but also delivers superior quantitative and qualitative results across various testing scenarios. This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving a way for personalized and more robust image restoration techniques.